Bob pays $10.00 for sunscreen without tax at walmart. He pays $10.83 with tax at walmart. Susan pays $5.00 for sunscreen at Target without tax. She pays $5.42 with tax at Target. What is the sale tax rate for both items?
Answer by checkley77(12844)

(10.83-10)/10=.83/10=.083=8.3% tax @ Walmart.
(5.42-5)/5=.42/5=.084=8.4% tax @ Target.